OpenStudy (anonymous):

What Are The Solutions of The System? y = x^2 + 3x - 4 y = 2x + 2 a. –1, 1) and (–7, –23) b. (–1, 1) and (7, 33) c. (–1, 33) and (7, 1) d. no solution

OpenStudy (anonymous):

From this set up you can substitute y y = x^2 + 3x - 4 y = 2x + 2 2x + 2 = x^2 + 3x - 4

OpenStudy (anonymous):

Put everything on one side so we can have it equal zero then use the quadratic formula

OpenStudy (anonymous):

Actually, you can just factor it after bringing everything to one side.

OpenStudy (anonymous):

Yeah i did that but then i get confused >.< ... I was leaning toward d as the answer but idk if i'm right

OpenStudy (anonymous):

Once you factor it, you get (x + 3)(x - 2) = 0 x = -3, 2 ummm....I don't know what happened in you problem...I'm sorry. :(

OpenStudy (anonymous):

Oh >.< I'm Not very good at algebra. But which is the solution A B C or D

OpenStudy (anonymous):

Honestly, I would say no solution because none of the other choices make sense...